# Dissecting the Hubble tension: Insights from a diverse set of Sound Horizon-free measurements

### v1

@@ IV.4 Bimodality in the One-Step Distribution
Although Figure 3 does not exhibit an obvious bimodal structure—the distribution of One-Step measurements appears reasonably concentrated around a single peak—a more detailed examination reveals a subtle but significant effect.
Upon closer inspection, we found that removing even two high-precision measurements (the Pogosian et al. [Pogosian2024] DESI and pre-DESI results) substantially alters the overall shape of the distribution. These measurements, with uncertainties of km s-1 Mpc-1, are among the most precise in our compilation. Due to their small uncertainties, they exert disproportionate influence on the inverse-variance weighted distribution, effectively pulling the combined probability density upward at their location (– km s-1 Mpc-1).
When these high-weight measurements are excluded, the inverse-variance weighted sum of Gaussians transforms into a distinctly bimodal curve shown in Figure 4. This suggests that the apparent unimodality in the raw distribution may be an artifact of a small number of highly precise measurements rather than a reflection of the true underlying structure.
We emphasize that this analysis should be interpreted with caution. Selectively removing measurements to reveal bimodality could be seen as a form of data manipulation. However, our purpose here is not to argue that the Pogosian et al. measurements are incorrect, but rather to illustrate that the statistical structure of the distribution is sensitive to a small number of high-precision anchors. A more rigorous approach would involve mixture model analysis (e.g., Gaussian Mixture Models) to formally test for bimodality, which we leave for future work.
@@ IV.5 Physical Interpretation of the Bimodality
Category 4 measurements anchored to early-universe physics (-based BAO)
@@
The High- Peak ( km s-1 Mpc-1): Populated by:
Specific measurements: Megamasers [Barua2025] and tailored EPM [Vogl2024]
These methods are characterized by their reliance on local geometric or kinematic physics.
This bifurcation pattern suggests a reformulation of the Hubble tension: the discrepancy may fundamentally be between model-dependent methods (which yield low ) and local astrophysics dependent or model-independent methods (which yield high ), rather than between early and late times.
@@ IV.6 The Role of High-Precision Anchors
Riess et al. [Riess2025] JWST Cepheids: km s-1 Mpc-1
Zhang et al. [Zhang2024] pairwise velocities: km s-1 Mpc-1
Vogl et al. [Vogl2024] tailored EPM: km s-1 Mpc-1
@@
Pogosian et al. [Pogosian2024] DESI BAO (no ): km s-1 Mpc-1
Piratova-Moreno et al. [PiratovMoreno2025] FRBs: km s-1 Mpc-1
Kelly et al. [Kelly2023] SN Refsdal: km s-1 Mpc-1
The weighted means of each category are heavily influenced by these anchor measurements. Consequently, resolving the Hubble tension requires either: (a) identifying systematic errors in one or more anchor measurements, or (b) explaining why different methodologies yield systematically different results.
@@ IV.7 Category 3 as the Critical Adjudication Ground
The “Pure Local One-Step” category (Category 3) serves as the critical test for distinguishing between the two interpretations of the Hubble tension. Currently, this category yields an intermediate value ( km s-1 Mpc-1), but it is internally heterogeneous:
@@
Low- subset: Zhang CC+HII (), Moresco CC (), Li lensing ()
@@
If Category 3 converges toward Category 1 (Distance Ladder): This would strongly imply that the CDM model assumption used in Category 2 is the source of the bias. Such convergence would suggest that the true late-time expansion history deviates from CDM, requiring new physics (e.g., dynamical dark energy, modified gravity, or a late-time gravitational transition)such as the CDM model [akarsu2024, Akarsu2025]
@@
Currently, the intermediate position of Category 3, combined with its internal heterogeneity, does not clearly favor either scenario. However, the existence of multiple independent model-independent measurements (EPM, Megamasers) that support the high- solution is suggestive.
@@ IV.8 Summary: The Model-Dependence Effect
The One-Step distribution shows evidence for bimodality when high-precision anchoring measurements are removed, with peaks corresponding to model-dependent () and model-independent () methods.
@@
Specific model-independent measurements (Vogl EPM) yield high values consistent with the Distance Ladder, suggesting the high value may reflect true local physics rather than distance ladder systematics.
@@
These findings support a recharacterization of the Hubble tension as primarily a “Distance Ladder vs. The Rest” crisis, with a secondary “Model-Dependent vs. Model-Independent” tension that may hold the key to its ultimate resolution.
@@ V Discussion and Conclusions
The “Hubble Tension” has long been framed as a cosmological crisis pitting the Early Universe (calibrated by the sound horizon) against the Late Universe (measured by the distance ladder). In this work, we have dissected this narrative by analyzing 83 measurements of that are strictly independent of the CMB Sound Horizon scale. Our findings suggest that the crisis is not necessarily a temporal one, but rather a methodological one. In this concluding section, we synthesize our results, discuss their implications for proposed solutions to the tension, and outline the observations needed to achieve resolution.
@@
As detailed in Section III, the Distance Ladder (Category 1) yields a robust high value of km s-1 Mpc-1. In stark contrast, when we combine all 54 independent measurements from Categories 2, 3, and 4—comprising Gravitational Waves, Strong Lensing, Cosmic Chronometers, Megamasers, FRBs, and sound-horizon-free CMB data—we obtain a combined value of km s-1 Mpc-1.
The resulting tension ( after correlation adjustment) between the Distance Ladder and this “All One-Step” ensemble exceeds the standard Planck vs. SH0ES discrepancy (). This indicates that the high value of is a unique feature of the three-rung calibration chain (involving Cepheids/TRGB/Miras SNe Ia) and is not supported by the broader consensus of Sound-Horizon-free cosmology.
@@
It is consistent with our 2024 analysis [Perivolaropoulos2024], despite the significant expansion of the dataset.
@@
V.2 Local Physics Does Not Imply High
A pervasive misconception in the field is that “local” measurements universally favor high values, thereby pointing to a breakdown of CDM at late times. Our analysis decisively refutes this assumption.
We found that 33 measurements utilizing exclusively local data () but assuming the validity of the CDM expansion history (Category 2) yield a weighted mean of km s-1 Mpc-1. This value is in perfect agreement with the Planck 2018 result ( km s-1 Mpc-1) [Planck2018]. Local probes such as Gravitational Wave standard sirens [Abbott2021], Strong Lensing time delays [Shajib2023, Kelly2023, Birrer2025], and FRBs [Wu2022, PiratovMoreno2025] do not intrinsically favor high . When analyzed under the standard model framework, local data supports the early-universe inference.
Furthermore, Category 4 demonstrates that even without the Sound Horizon , CMB data constraints (driven by the matter-radiation equality scale ) favor a low km s-1 Mpc-1 [Pogosian2024, Bahr-Kalus2025].
Thus, neither “Local” nor “Sound Horizon free” is synonymous with “High .” The high value may be specific to the Distance Ladder methodology.
V.3 Internal Tension and the Shape of
While the external tension with the Distance Ladder dominates the landscape, we have identified a secondary internal tension among the One-Step measurements that may hold the key to understanding the underlying physics.
Measurements that assume the CDM shape of (Category 2) yield km s-1 Mpc-1. However, measurements that are free from this shape assumption (Category 3) yield a higher intermediate value of km s-1 Mpc-1. This km s-1 Mpc-1 shift accounts for more than half of the total Hubble tension.
This systematic offset has two possible interpretations:
1.
CDM is correct: Model-independent methods (Category 3) suffer from larger uncontrolled systematics that bias high.
2.
CDM is incomplete: The true expansion history deviates from the CDM prediction at , and forcing a CDM fit biases low.
The fact that specific model-independent measurements—particularly the Vogl et al. [Vogl2024] tailored EPM ()— yield values consistent with the Distance Ladder is intriguing. This method is independent of both the Cepheid/SNe Ia calibration chain and cosmological model assumptions, yet it supports the high- solution. This pattern is more naturally explained by interpretation (2) than by a conspiracy of unrelated systematics.
The persistence of this internal tension reinforces earlier findings that the CDM model may face challenges beyond the Hubble tension alone. A comprehensive review of these challenges [Perivolaropoulos2022review] identifies correlated anomalies that may point toward a common origin in modified late-time physics.
V.4 Implications for Proposed Solutions
@@
V.4.1 Early Dark Energy and Sound Horizon Modifications
Models that resolve the tension by reducing the Sound Horizon at recombination (e.g., Early Dark Energy [Knox2020, Mirpoorian2025], extra relativistic species, early modified gravity) predict that Sound-Horizon-independent measurements should align with the Distance Ladder. Our analysis shows the opposite: Sound-Horizon-free measurements (Categories 2, 3, and 4) predominantly align with Planck, not with SH0ES.
@@ V.4.1 Early Dark Energy and Sound Horizon Modifications
The failure of most Sound-Horizon-independent probes to support the high- solution disfavors pure early-universe modifications as the complete explanation.
V.4.2 Late-Time Modifications
@@
Recent DESI results [DESI2024, Giare:2025pzu, Gao:2024ily] suggesting evolving dark energy (, ) may provide partial support, but the statistical significance remains modest and the robustness across different parameterizations has been questioned [Colgain:2024mtg]. Theoretical frameworks for such behavior include models where the effective gravitational constant undergoes a rapid transition at late times, which can arise naturally in scalar-tensor theories. Recent re-examinations of such modified local physics scenarios using SH0ES data have refined the constraints on these models. A particularly intriguing perspective recasts the Hubble tension as an crisis. Theodoropoulos & Perivolaropoulos demonstrated that if the intrinsic luminosity of SNe Ia in the Hubble flow differs systematically from that in calibrator galaxies, the inferred would be biased without requiring modifications to fundamental cosmology [Marra2021, Perivolaropoulos:2025gzo, Theodoropoulos2021] .
V.4.3 Distance Ladder Systematics
The isolation of the Distance Ladder as a outlier against 54 independent measurements points toward unrecognized systematics within the calibration chain. Potential systematic effects include:
@@ V.4.3 Distance Ladder Systematics
SNe Ia systematics: Environmental dependencies (host galaxy mass, local star formation), evolution with redshift, or standardization procedure biases.
@@
Multiple independent calibrators (Cepheids, TRGB, JAGB, Miras, SBF) yield consistent high- values.
JWST observations have largely confirmed HST Cepheid photometry [Riess2025].
Model-independent methods like the tailored EPM [Vogl2024] support the high- solution.
V.4.4 Hybrid Scenarios
@@
V.5 Falsifiability and Future Tests
@@
V.5.1 What Would Confirm Distance Ladder Systematics?
@@
V.5.2 What Would Confirm New Physics?
@@
Multiple independent model-independent methods (EPM, Megamasers, future techniques) consistently yield km s-1 Mpc-1.
@@ V.6 Observational Priorities
Expand Category 3 (Pure Local One-Step): This category currently contains only 15 measurements with significant internal heterogeneity. Additional model-independent measurements are crucial for determining whether the intermediate km s-1 Mpc-1 persists or converges toward one of the two solutions.
@@
Improve tailored EPM precision: The Vogl et al. [Vogl2024] result is currently the strongest model-independent support for the high- solution outside the traditional Distance Ladder. Larger SNe II samples observed with JWST spectroscopy could significantly reduce uncertainties.
@@
Accumulate GW standard sirens: The LIGO-Virgo-KAGRA O4 and O5 [LIGOScientific:2025jau] runs will significantly increase the sample of both bright and dark sirens. Crucially, these should be analyzed both with and without CDM assumptions to test for model-dependent biases.
@@
Model-independent Strong Lensing: Future time-delay measurements should emphasize cosmographic or model-independent analyses alongside standard CDM fits.
@@
V.7 Concluding Remarks
The Hubble tension remains one of the most significant challenges to the standard cosmological model. Our analysis of 83 Sound-Horizon-independent measurements reveals that the tension is more accurately characterized as a “Distance Ladder vs. The Rest” crisis rather than an “Early vs. Late Universe” discrepancy.
@@ V.7 Concluding Remarks
The Distance Ladder yields km s-1 Mpc-1, while all other Sound-Horizon-free methods combined yield km s-1 Mpc-1, a ( correlation-adjusted) tension.
@@
A internal tension exists between CDM-dependent (Category 2) and model-independent (Category 3) local measurements, with the CDM assumption reducing inferred by km s-1 Mpc-1.
@@
Specific model-independent measurements (tailored EPM,) support the high- solution, complicating the pure “systematics” interpretation.
@@
Whatever the ultimate explanation, the Hubble tension has already achieved a valuable scientific outcome: it has motivated an unprecedented scrutiny of both our measurement techniques and our theoretical assumptions, pushing cosmology toward greater rigor and potentially toward new discoveries.Recent results from the Atacama Cosmology Telescope DR6 provide additional CMB constraints that will be crucial for further testing the robustness of our conclusions [AtacamaCosmologyTelescope:2025blo, AtacamaCosmologyTelescope:2025nti]

### v2

@@ IV.4 Bimodality in the Distance Ladder Independent/Sound Horizon Free (DLI–SHF) Distribution
Although Figure 3 does not exhibit an obvious bimodal structure—the distribution of DLI–SHF measurements appears reasonably concentrated around a single peak—a more detailed examination reveals a subtle but significant effect.
To further explore the underlying structure of the DLI–SHF measurements, demonstrated also in figure 2, we employ a weighted Kernel Density Estimation (KDE) approach. Unlike a simple inverse-variance weighted sum of Gaussians, where each measurement contributes a Gaussian with width equal to its uncertainty , the KDE method assigns a fixed bandwidth to all measurements while weighting their contributions according to their precision. The weighted KDE is defined as:
where is the measured value, is the bandwidth (kernel width), and are the normalized inverse-variance weights:
This approach ensures that high-precision measurements contribute proportionally through their weights, but do not dominate the distribution through both weight and narrow kernel width simultaneously—a problem that arises with the standard inverse-variance weighted sum of Gaussians.
For this analysis, we exclude Category 4 measurements (CMB Sound Horizon Free), as these rely on CMB-derived data that provide exceptionally high precision. The three most precise Category 4 measurements—Pogosian et al. [144, 145] (DESI: km s-1 Mpc-1; pre-DESI: km s-1 Mpc-1) and Garcia et al. [70] ( km s-1 Mpc-1)—would alone account for approximately of the total inverse-variance weight, clustering near km s-1 Mpc-1 and obscuring any underlying structure in the remaining measurements.
The bandwidth is a critical parameter controlling the smoothness of the KDE. Silverman’s rule of thumb [167], which minimizes the mean integrated squared error assuming a unimodal Gaussian distribution, yields km s-1 Mpc-1 for our data. However, this criterion is known to oversmooth multimodal distributions. Given the a priori distinction between Category 2 (CDM-dependent) and Category 3 (model-independent) measurements, a smaller bandwidth is justified to reveal potential substructure.
We find that bimodal structure is clearly visible for bandwidths in the range km s-1 Mpc-1. For , additional peaks emerge due to over-resolution of statistical fluctuations; for , the two peaks merge into a single mode with a shoulder. We adopt km s-1 Mpc-1, an intermediate value within the robust range, to balance resolution and smoothing.
The resulting distribution, shown in Figure 4, exhibits two distinct peaks corresponding to the weighted means of Category 2 ( km s-1 Mpc-1) and Category 3 ( km s-1 Mpc-1) measurements, with a separation of km s-1 Mpc-1. This bimodal structure suggests a systematic difference between CDM-dependent and cosmological-model-independent not Distance Ladder methods that warrants further investigation.
@@ IV.5 Physical Interpretation of the Bimodality
This bifurcation pattern suggests a reformulation of the Hubble tension: the discrepancy may fundamentally be between model-dependent methods (which yield low ) and local astrophysics dependent or cosmological-model-independent methods (which yield high ), rather than between early and late times.
@@ IV.6 The Role of High-Precision Anchors
Multi-parameter transition analyses of SH0ES data have explored how calibrator parameters may vary across the distance ladder, potentially contributing to the observed bifurcation. Evidence for possible transitions in Cepheid-SNe Ia calibrator parameters has been previously identified in the SH0ES sample [129, 164]. Recent work has also highlighted the subtle statistical effects in Distance Ladder analyses, including distance priors and selection effects, that could impact such comparisons [46]. The weighted means of each category are heavily influenced by these anchor measurements. Consequently, resolving the Hubble tension requires either: (a) identifying systematic errors in one or more anchor measurements, or (b) explaining why different methodologies yield systematically different results.
@@ IV.7 Category 3 as the Critical Adjudication Ground
The “Pure Local category (Category 3) serves as the critical test for distinguishing between the two interpretations of the Hubble tension. Currently, this category yields an intermediate value ( km s-1 Mpc-1), but it is internally heterogeneous:
@@
Low- subset: Zhang CC+HII (), Li lensing ()
@@
If Category 3 converges toward Category 1 (Distance Ladder): This would strongly imply that the CDM model assumption used in Category 2 is the source of the bias. Such convergence would suggest that the true late-time expansion history deviates from CDM, requiring new physics (e.g., dynamical dark energy, modified gravity, or a late-time gravitational transition)such as the CDM model [6, 5]
@@
Currently, the intermediate position of Category 3, combined with its internal heterogeneity, does not clearly favor either scenario. However, the existence of multiple independent cosmological-model-independent measurements (EPM, Megamasers) that support the high- solution is suggestive.
@@ IV.8 Summary: The Model-Dependence Effect
The distribution shows evidence for bimodality with peaks corresponding to model-dependent () and cosmological-model-independent () methods.
@@
Specific cosmological-model-independent measurements (Vogl EPM) yield high values consistent with the Distance Ladder, suggesting the high value may reflect true local physics rather than distance ladder systematics.
@@
These findings support a recharacterization of the Hubble tension as primarily a “Distance Ladder vs. The Rest” crisis, with a secondary “Model-Dependent vs. Cosmological-model-Independent” tension that may hold the key to its ultimate resolution.
@@ V Discussion and Conclusions
The “Hubble Tension” has long been framed as a cosmological crisis pitting the Early Universe (calibrated by the Sound Horizon) against the Late Universe (measured by the distance ladder). In this work, we have dissected this narrative by analyzing 88 measurements of that are strictly independent of the CMB Sound Horizon scale. Our findings suggest that the crisis is not necessarily a temporal one, but rather a methodological one. In this concluding section, we synthesize our results, discuss their implications for proposed solutions to the tension, and outline the observations needed to achieve resolution.
@@
As detailed in Section III, the Distance Ladder (Category 1) yields a robust high value of km s-1 Mpc-1. In stark contrast, when we combine all 58 independent measurements from Categories 2, 3, and 4—comprising Gravitational Waves, Strong Lensing, Cosmic Chronometers, Megamasers, FRBs, and Sound-Horizon-Free CMB data—we obtain a combined value of km s-1 Mpc-1.
The resulting tension ( after correlation adjustment) between the Distance Ladder and the Distance Ladder Independent/Sound Horizon Free (DLI–SHF) ensemble exceeds the standard Planck vs. SH0ES discrepancy (). This indicates that the high value of is a unique feature of the three-rung calibration chain (involving Cepheids/TRGB/Miras SNe Ia) and is not supported by the broader consensus of Sound-Horizon-Free cosmology.
@@
It is consistent with our 2024 analysis [134], despite the significant expansion of the sample.
@@
V.2 Robustness of Results
To assess the sensitivity of our results to methodological choices in measurement categorization, we performed a comprehensive robustness analysis111We thank Dan Scolnic for suggesting the examination of the categorization sensitivity for these specific cases to test the robustness of the internal tension.. We considered a more conservative approach to categorization by: (i) adopting alternative reported values for measurements with multiple H0 estimates (Liu et al.: 59.1 67.5 km s-1 Mpc-1; Dominguez et al.: 61.9 65.6 km s-1 Mpc-1), (ii) removing potentially redundant measurements from the same observational data (Kelly/Li Refsdal analyses), and (iii) excluding measurements whose categorization could be debated (Zhang et al. 2022 cosmic chronometer analysis). Table 8 summarizes the impact of these changes on our tension estimates.
| Scenario | Cat2 | Cat3 | Tension | Tension |
| (DL-LI) | (C2-C3) | |||
| Baseline (this work) | ||||
| Liu: 59.167.5 | ||||
| Dominguez: 61.965.6 | ||||
| Remove Kelly (Refsdal) | ||||
| Remove Li (Refsdal) | ||||
| Remove Zhang 2022 | ||||
| Liu + Dominguez | ||||
| All changes (keep Li) | ||||
| All changes (keep Kelly) |
The analysis reveals that our main result—the significant tension between Distance Ladder and all others measurements—is robust to all tested variations in categorization methodology. Even under the most conservative scenario combining all alternative choices, the main tension remains at , well above the threshold for statistical significance. No individual change affects the main tension by more than .
In contrast, the internal tension between Category 2 (CDM-dependent) and Category 3 (cosmological-model-independent) measurements is more sensitive to these methodological choices, varying from to depending on the specific categorization adopted.
V.3 Implications for Proposed Solutions
@@
V.3.1 Early Dark Energy and Sound Horizon Modifications
Models that resolve the tension by reducing the Sound Horizon at recombination (e.g., Early Dark Energy [91, 114, 173], extra relativistic species, early modified gravity) predict that Sound-Horizon-Free measurements should align with the Distance Ladder. Our analysis shows the opposite: Distance Ladder Independent/Sound Horizon Free (DLI–SHF) (Categories 2, 3, and 4) yields a value closer to Planck, not to SH0ES.
@@ V.3.1 Early Dark Energy and Sound Horizon Modifications
The failure of most Sound-Horizon-Free probes to support the high- solution disfavors pure early-universe modifications as the complete explanation.
V.3.2 Late-Time Modifications
@@
Recent DESI results [43, 72, 68] suggesting evolving dark energy (, ) provide evidence for deviations from CDM at late times. However, due to the strong anti-correlation between and [40], such models with actually predict lower values, potentially exacerbating rather than alleviating the tension with the Distance Ladder. Moreover, the statistical significance of the DESI signal remains modest (–) and its robustness across different parameterizations has been questioned [39]. Theoretical frameworks for such behavior include models where the effective gravitational constant undergoes a rapid transition at late times, which can arise naturally in scalar-tensor theories. Recent re-examinations of such modified local physics scenarios using SH0ES data have refined the constraints on these models. A particularly intriguing perspective recasts the Hubble tension as an crisis. Theodoropoulos & Perivolaropoulos demonstrated that if the intrinsic luminosity of SNe Ia in the Hubble flow differs systematically from that in calibrator galaxies, the inferred would be biased without requiring modifications to fundamental cosmology [112, 128, 177].
V.3.3 Distance Ladder Systematics
The isolation of the Distance Ladder as a outlier against 58 independent measurements points toward unrecognized systematics within the calibration chain. Potential systematic effects include:
@@ V.3.3 Distance Ladder Systematics
V.3.4 Hybrid Scenarios
@@
V.4 Falsifiability and Future Tests
@@
V.4.1 What Would Confirm Distance Ladder Systematics?
@@
V.4.2 What Would Confirm New Physics?
@@
Multiple independent cosmological-model-independent methods (EPM, Megamasers, future techniques) consistently yield km s-1 Mpc-1.
@@ V.5 Observational Priorities
Expand Category 3 (Pure Local): This category currently contains only 16 measurements with significant internal heterogeneity. Additional cosmological-model-independent measurements are crucial for determining whether the intermediate km s-1 Mpc-1 persists or converges toward one of the two solutions.
@@
Improve tailored EPM precision: The Vogl et al. [182] result is currently the strongest cosmological-model-independent support for the high- solution outside the traditional Distance Ladder. Larger SNe II samples observed with JWST spectroscopy could significantly reduce uncertainties.
@@
Accumulate GW standard sirens: The LIGO-Virgo-KAGRA O4 and O5 [1] runs will significantly increase the sample of both bright and dark sirens. Crucially, these should be analyzed both with and without CDM assumptions to test for model-dependent biases.
@@
Cosmological-model-independent Strong Lensing: Future time-delay measurements should emphasize cosmographic or cosmological-model-independent analyses alongside standard CDM fits.
@@
V.6 Concluding Remarks
The Hubble tension remains one of the most significant challenges to the standard cosmological model. Our analysis of 88 Sound-Horizon-Free measurements reveals that the tension is more accurately characterized as a “Distance Ladder vs. The Rest” crisis rather than an “Early vs. Late Universe” discrepancy.
@@ V.6 Concluding Remarks
The Distance Ladder yields km s-1 Mpc-1, while all other Distance Ladder Independent/Sound Horizon Free(DLI–SHF) methods combined yield km s-1 Mpc-1, a ( correlation-adjusted) tension.
@@
A internal tension exists between CDM-dependent (Category 2) and cosmological-model-independent (Category 3) local measurements, with the CDM assumption reducing inferred by km s-1 Mpc-1.
@@
Specific cosmological-model-independent measurements (e.g., tailored EPM) support the high- solution, complicating the pure “systematics” interpretation.
@@
Whatever the ultimate explanation, the Hubble tension has already achieved a valuable scientific outcome: it has motivated an unprecedented scrutiny of both our measurement techniques and our theoretical assumptions, pushing cosmology toward greater rigor and potentially toward new discoveries.Recent results from the Atacama Cosmology Telescope DR6 provide additional CMB constraints that will be crucial for further testing the robustness of our conclusions [108, 28]